OpenStudy (anonymous):

Solve the equation below for y. 8x - 2y = 24 A. y= 4x - 12 B. y= 12 - 4x C. y = 8x - 24 D.y = 4x - 8

OpenStudy (anonymous):

Help i need to be guided towards the right way please help me !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

8x - 2y = 24 Add -8x on both sides

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

hold on a minute

OpenStudy (anonymous):

what you get is -2y=24-8x

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Now, divide both sides by -2

OpenStudy (anonymous):

i got y=12-4x and that is my answer thank you

OpenStudy (anonymous):

No.....that dont look correct

OpenStudy (anonymous):

y = 12 - 4x

OpenStudy (anonymous):

-2y=24-8x y=24/(-2) -8x/(-2) y=-12+4x y=4x-12

OpenStudy (anonymous):

u missed the -ve sign

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

great

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome